40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  TCS elog  Not logged in ELOG logo
Message ID: 76     Entry time: Mon Jul 26 09:42:30 2010
Author: Aidan 
Type: Computing 
Category: SLED 
Subject: Long term test on SLED started - Day 0 

 I set up the SLED to test its long term performance. The test began, after a couple of false starts, around 9:15AM this morning.

The output of the fiber-optic patch cord attached to the SLED is illuminating a photo-detector. The zero-level on the PD was 72.7mV (with the lights on). Once the PD was turned on the output was ~5.50 +/- 0.01V. This is with roughly 900uW exiting the SLED.

The instructions from Superlum suggest limiting the amount of power coupled back into the fiber to less than 3%. With the current setup, the fiber is approximately 2" from the photodetector. What is the power coupled back into the fiber?

Assume a worst case of 100% of the light reflected from the PD, the wavelength is 830nm and a waist size of about 6um radius at the output of the fiber. The beam size at 4" (from the fiber output to the PD and back again) or ~100mm from the fiber is about 4.4mm radius. Therefore about (6um/4.4mm)^2 or ~2ppm will be coupled back into the fiber. This is sufficiently small.

The attached plots from dataviewer show measurements from the SLED (on-board photodetector, on-board temperature sensor, current setpoint, current limit, current to diode) over the last 15 hours.

Attachment 1: SLED_superlum_long_term_test_0001A.pdf  60 kB  Uploaded Mon Jul 26 11:05:35 2010  | Hide | Hide all
Attachment 2: SLED_superlum_long_term_test_0001B.pdf  59 kB  Uploaded Mon Jul 26 11:05:50 2010  | Hide | Hide all
ELOG V3.1.3-